OpenStudy (anonymous):

6 orders of Triple integration 0

myininaya (myininaya):

is there suppose to be a function?

myininaya (myininaya):

\[\int\limits_{?}^{?}\int\limits_{?}^{?}\int\limits_{?}^{?}f(x,y,z) dx dy dz\] ?

myininaya (myininaya):

i'm not sure what the 6 orders of part means

OpenStudy (agreene):

myininaya the inequalities are those limits you have as ?

myininaya (myininaya):

yes i know but what is f(x,y,z)?

myininaya (myininaya):

i just didn't write the limits

OpenStudy (anonymous):

there is no function

OpenStudy (anonymous):

it just wants you to find the limits

myininaya (myininaya):

i don't get it the limits are given

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\int\limits_{?}^{?}\int\limits_{?}^{?}\] for all 6 orders though... so dydxdz, dxdydz, dzdxdy

OpenStudy (anonymous):